OpenStudy (anonymous):

the first number is eight more than the second number. three times the second number plus twice the first number is equal to 26. find the numbers

OpenStudy (anonymous):

let first no. = x and secnd = y. x = y + 8 2x + 3y = 26..........a 2x = 2y + 16 2x -2y = 16.............b subtract b from a: 5y = 10 y = 2 and x = 2+8 = 10

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

what is the . . . . a for ?

OpenStudy (anonymous):

i'm just identifying 2 of the equations by a and b instead of repeating the whole equations. i could have called them 1 and 2 or any other 2 nunbers or letters

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

yw